Ivan (John) Zacharuk Katherine Antoniuk  Michael Antoniuk 1872-1952  Maria Toffin 1878-1972  Fedun Antoniuk 1876?-?  ?  William Turyk 1872-1941 Donna Antoniuk 1879-1974  Alex Yurechuk 1878-1963  Mary Antoniuk 1882-1978  John (Swidzinski) Swedinski 1886-1976  Olena (Ellen) Antoniuk 1884-1982  Nicholas Antoniuk 1890-1949 Katherine Dubetz 1896-1964 m  m  m 1898 m  m  m 1913  m Alex Wintonyk ?-1969  Maria Zacharuk ?-1985  ? Derbowka Donna Zacharuk Stephen Zacharuk  ?  Frank (Martiniuk) Martin  Ann Zacharuk  ?  Florence Zacharuk  ? Acktemchuk Nellie Zacharuk  ? Wyllchuk  Nita Zacharuk Michael Zacharuk  ?  Nicholas Zacharuk ?-1993  ?  Fred Zacharuk  ? m  m  m  m  m  m  m  m  m  m Anne Wintonyk 1915-  Helene Wintonyk 1917-  Nicholas Wintonyk 1919-  Marie Wonnetoway 1922-  Anita Wintonyk 1921- m Wylsen Wintonyk  Terry Wintonyk  Diana Wintonyk  Roman Antoniuk 1901-1960  Mary Sawchuk 1906-1978  Nicholas Warzin 1893-? Katherine Antoniuk 1904-1993  Nicholas Antoniuk 1909-1989  William Antoniuk 1911-1991  Annie Sawchuk 1911-  Donna Antoniuk 1913-  Stephen Antoniuk 1915-1977  Mary Lobay 1919-  Stephen Zytaruk 1917-1988 Lillian Antoniuk 1917-  Michael Antoniuk 1919-  Kathleen Nylen 1918-  Ernest Hickmore 1924-  Mary Antoniuk 1921- m  m  m  m 1942  m 1942  m  m William Michael Antoniuk 1930-1982  Violet Gutek 1933-1974 Joanne (Barber) Gillson 1926-  Douglas Eugene Thompson 1928-1985  Ellen Marie Antoniuk 1932-  Robert Michael Elaschuk 1945-  Elizabeth Anne Antoniuk 1945- m  m  m 1950  m 1964 William Michael Antoniuk 1956-  Stephen Roman Antoniuk 1966-  Malcolm Henderson 1952-  Shelley Anne Thompson 1954-  Douglas Roman Thompson 1956-  Marliss Gibson 1957-  Wendy Brandvold 1960- m 1972  m 1977  m 1984 Ryan Douglas Thompson 1975-  Jaeson Robert Thompson 1977-  Collin Brent Thompson 1980-  Jade Joanne Thompson 1978-  Travis Douglas Thompson 1984-  Robinn Kary Elaschuk 1977-  Mary Warzin  Michael Warzin 1925-  Elizabeth (Elsie) Wakaruk 1933-  John Lukien  Mary Warzin 1928-  Harry Warzin 1930-  June Generoux  William Warzin 1933-  Phyllis Murik 1937- m 1953  m 1947  m  m 1955 Darlene Warzin 1957-  ?  Elaine Warzin 1964-  Audrey Warzin 1966- m  Richard Lukien 1949-19?? Murray Scott  Shirley Lukien 1952- m 1971 Tracy Scott 1972-  Valerie Scott 1973-  Dennis Warzin 1954-  Donna Mattock  Ronald Warzin 1957-  Glenna Young  ?  Joanne Warzin 1965-  Alexander Warzin 1978- m 1976  m 1978  m  Douglas Henderson Dwayne Braemer  Marian Warzin  Elmer Warzin 1957-  ?  Nancy Warzin 1961-1994  Taras Warzin 1969-  ? m  m  m  m Paul Warzin 1972- Randy Henderson 1975-  2 or 3 marriages 3 children  Child  Child  Martin William Antoniuk 1937-  Eunice Ratsoy 1937-1970 Majorie (Bonnie) Shaw  Leslie Michael Antoniuk 1942-  Mary Norman  Diane Repyen Heacock 1946- m  m 1968  m  m 1976 Ward William Antoniuk 1961-  Karen Dobey 1961-  Nathan Wade Antoniuk 1962-  Lonnie Shulko m 1992  m Kennedy Anne Antoniuk 1994-  Richard Antoniuk 1943-  Wendy ?  Randolph Antoniuk 1947-  Jane Adams 1947-  Rodney Antoniuk Maureen (Moe) Riley 1949- m  m 1971  m 1975 Lucas Antoniuk 1976-  Child  Child  Child  Barry Zytaruk 1944-  Lee Wirun  Murray Zytaruk 1952-  Muriel Keller m 1972  m 1972 James Zytaruk  William Zytaruk  Gregory Barnes 1954-  Barbara Lynn Antoniuk 1953-  Howard Michael Antoniuk 1956-  Joan Dzigolick m  m Heather Marie Barnes 1980-  Brandi Michelle Antoniuk 1986-  Dmytro Fodchuk 1930-1991 Sophia Antoniuk 1900-1993  Mary Antoniuk 1903-1971? Nicholas Antoniuk ?-1981?  Stephen Antoniuk  Mary ? m  m William Fodchuk 1930-  June Helene McLennan 1931- m 1951 Richard William Fodchuk 1951-  Lena Roppo 1964-  Mark (Fodchuk) McLennan 1953-  Marian Rindero  Tara Rhien (Hoover) Annalee Jil Fodchuk 1956-  Michael Geordie Fodchuk 1960-  Jason McLennan Fodchuk 1966- m 1984 Joshua Shane Fodchuk 1988-  Ryan William Fodchuk 1994-  Layla Rindero 1972-  William Clayton McLennan 1989-  Katherine Antoniuk  Stephen Turyk 1904-1978 Katherine Dubetz 1896-1964 Michael Turyk 1912-  Helen Sawchuk 1917- m  m 1941 David Lewison 1940-  Olive Julie Turyk 1940-  Peter Paul Turyk 1942-  Judy Durward 1946-  Joyce Victoria Turyk 1946-1947 Dennis Michael Turyk 1949-  Kirk Whitfield 1952-  Terri Anne Turyk 1956- m 1970  m 1968  m 1975 Justin David Lewison 1978-  Jason Michael Turyk 1970-  Belinda ?  Matthew Robert Turyk 1973-  Anne Melissa Turyk 1978- m Braedon Mitchell Turyk 1994-  Stephen Yurechuk 1902-1971 Katherine Kolody 1908-  Stephen Michajluk 1894-1977 Donna Yurechuk 1905-1991  William Finnie  Ellen Yurechuk 1907-  William Kassian 1905-  Anna Yurechuk 1910-1989  William Yurechuk 1913-1989  Olga Maksymic 1923-1976  David Bengtson 1914-  Lillian Yurechuk 1915-1966  William (Formanski) Forman 1913-1986 William (Simkovich) Simcoe  Mary Yurechuk 1917-1972  Alexander Yurechuk 1924-  Gladys Lundeby 1924- m  m 1927  m  m  m  m 1941?  m  m  m Andrew (Yurechuk) Stephen 1927-  Anne ?  Elizabeth Josephine Summers Terrance (Yurechuk) Stephen 1930-  Stella Piasecki 1930-  Shirley A. Madison 1934- m  m 1957  m 1951  m 1981 Jane Elizabeth Stephen 1958-  Boyd Andrew  Ross Terrence Stephen 1956-  Dianne Schmitz  Rodney A. Stephen 1959-  Cheryl Caper  Stacey-Jo Stephen 1963- m 1982  m Tyler Stephen  Scott Stephen  Heather Stephen 1991-  Ernest George Lake 1928-  Mary Pauline Michajluk 1931-  William Stephen Michajluk 1936-  Ann Elizabeth Dagg m 1952  m 1964 Randall George Lake 1954-  Robert Murray Lake 1957-  Ronald John Willis 1960-  Debra Marilyn Lake 1959- m 1985 Shannon Marie Willis 1988-  Alanna Michelle Willis 1989-  Shawn Douglas Michajluk 1967-  Glen Robertson  Lisa Suzanne Michajluk 1970-  Peter William Michajluk 1972- m 1996  Terence Kassian 1933-  Wilma Steckel  Leonard Kassian  Mabel Yewchuk m  m Robert Kassian  Wayne Kassian  Carol Kassian  Brian Kassian  Lisa Kassian  William Yurechuk 1942-  Gladys Skordbahach  Douglas Kempf  Elizabeth Yurechuk 1944-  Douglas Pushor  Patricia Yurechuk 1955-  Terrence Fosvelt  Sandra Yurechuk 1961- m  m  m  m Carrie Yurechuk Stephen Yurechuk  Stacey Yurechuk Michael Yurechuk  Fred Kempf  Jo Kempf  Sheila Pushor  Bonnie Pushor  Colleen Pushor  Brandi Fosvelt  Dustin Fosvelt  Roald Dyrland 1939-  Judith Bengtson 1943-  Clifton Bengtson 1944-  Kay Luna m 1963  m  Gerald Forman 1937-  Rose ?  Vivian ?  Robert Forman 1939-  Patricia Mallick 1941-1993  Kenneth Forman 1942-1992 Sherry ? m  m  m  m Son  Son  Son  Daughter  Darren Forman  Kathy Forman  Son  Gerald Vetsch 1951-  Shirley Yurechuk 1951-  Gerald Kawchuk 1955-  Laurie Yurechuk 1955- m  m Shauna Vetsch 1974-  Ryan Vetsch 1976-  Brandon Vetsch 1978-  Craig Vetsch 1981-  Bradley Kawchuk 1983-  Jennifer Kawchuk 1987-  Michael (Swedinski) Swiss 1914-1994  Ina Thom 1920-1987  Nicholas Swedinski 1916-  Walter Swedinski 1918-1985  Anne Manweiler 1926-  Michael Burtnik 1917-  Anne Swedinski 1919-  Kenneth Myer 1925-  Stephanie Swedinski 1921-  Harry Kerry 1922-1993  Emily Swedinski 1923-  John Swedinski 1927-1994  Eileen Tripp 1926-1993 m  m  m  m  m  m 1950 Michael Hatzinikolas 1943-  Linda Swiss 1948-  James Swiss 1953-  Wendy Rogowski 1963- m  m Andrea Hatzinikolas 1977-  Stacy Hatzinikolas 1980-  Christina Hatzinikolas 1983-  Alexander Hatzinikolas 1984-  Sara Swiss  Brett Swiss 1984-  Marlee Swiss 1988- Randy Swedinski 1949-  Richard Burtnik 1943-  Natalie Shupenia 1949-  Walter Mis 1941-  Joan Burtnik 1945- m  m Lynnel Burtnik 1972-  Michael Burtnik 1974-  Peter Mis 1970-  Dominick Mis 1971-  Matthew Mis 1975-  Cory Myer 1954-  Terry Knipelberg 1953-  Darcy Myer 1958- m Christopher Myer 1984-  Jennifer Myer 1988-  Normin Simpson 1944-  Lauren Kerry 1951-1992  Craig Kerry 1953-  Celeste Shallian 1959- m  m Cole Kerry 1992-  Arthur Swedinski 1949-  Robert Swedinski 1951-  Sidney Swedinski 1953-  Robert Nagy  Eileen Swedinski 1955-  Michael Swedinski 1957-  Patricia ?  Kevin Johnson Margaret Swedinski 1961- m  m  m Shelley-Mae Nagy 1977-  Clayton Nagy 1984-  Darryl Swedinski  Allan Swedinski Michelle Swedinski 1992-  Adam Johnson 1988-  Aimee Johnson 1990-  Amanda Johnson 1992-

1